OpenStudy (anonymous):

Using Descartes? rule of signs, how many real roots does f(x) = x^3 + x^2 - 3x + 2 contain?

OpenStudy (anonymous):

My teacher didnt explain it well and i only have one example in my notes

OpenStudy (anonymous):

no positive real roots, no negative real roots two or zero positive real roots, one or zero negative real roots one or zero positive real roots, two or zero negative real roots two or zero positive real roots, two or zero negative real roots ??????

OpenStudy (anonymous):

I think it may be the last one? can someone help me understand this?

OpenStudy (whpalmer4):

Check out my post in your last question for an example of how to use the rule of signs...

OpenStudy (whpalmer4):

We have a highest power of \(x^3\) here, so there are 3 roots. We know that because this polynomial only has real coefficients, any complex roots come in pairs. That means our options are 1 real, 2 complex, or 3 real. Any answer that would require 1 or 3 complex roots isn't going work here.

OpenStudy (whpalmer4):

Hopefully you can wrap your head around why that means the last one cannot be correct.

OpenStudy (anonymous):

Okay yes i understnad

OpenStudy (whpalmer4):

doing the procedure here as I try to keep my eyelids propped open: x^3+x^2-3x+2 has 2 sign changes, so 2 or 0 positive real roots -x^3+x^2+3x+2 has 1 sign change, so 1 negative real root our only possible combinations that add to 3 roots are 2 positive, 1 negative, 0 complex or 0 positive, 1 negative, 2 complex

OpenStudy (whpalmer4):

in fact, the truth of the matter is that we have 1 negative at \( x = -2.51155\) and 2 complex at \( (x = 0.755774 \pm 0.474477i\)\)

OpenStudy (anonymous):

little confusing lol i get how to do the positive real roots but not the negative? How do you know which signs to change

OpenStudy (whpalmer4):

What you do for the negative is rewrite f(x) as f(-x), but to do that, all you have to do is go through the equation and change the signs of the terms where x is raised to an odd power (1,3,5...)

OpenStudy (whpalmer4):

That's because \((-x)^{2n} = x^{2n}\) but \((-x)^{2n+1} = -(x^{2n+1})\)

OpenStudy (anonymous):

Ohhh so dont change the even powers?

OpenStudy (whpalmer4):

even powers are unchanged when we substitute \(-x\) for \(x\), right? \[f(x) = x^2\]\[f(-x) = (-x)^2 = (-x)*(-x) = x^2\]

OpenStudy (anonymous):

Yeahh it all makes sense now! haha

OpenStudy (anonymous):

That has litterally been confusing me for hours i didnt realize how simple it was till you explained it lol

OpenStudy (whpalmer4):

It's one of those things that makes you say "yeah, uh huh, I understand that" as you watch someone do it, but then when you have to do it yourself the first few times, that "I understand" turns into "wait, WTF?" :-)

OpenStudy (anonymous):

Hahaha really though!!

OpenStudy (anonymous):

So for f(x) = x^4 + 3x^2 + 2 this would have no positive real roots, no negative real roots, correct?

OpenStudy (whpalmer4):

f(x): + + + no sign changes, no positive real roots f(-x): + + + no sign changes, no negative real roots must have 4 roots, so we have 2 pairs of complex conjugate roots

OpenStudy (whpalmer4):

Roots for this are \[x = \pm i, \,\,x = \pm i\sqrt{2}\]

OpenStudy (anonymous):

Okay kind of following you. so my answer is correct though?

OpenStudy (whpalmer4):

Yes, you were correct, as far as you went.

OpenStudy (whpalmer4):

I just was showing you my shorthand notation: \[x^4+3x+2\]becomes + + + because that is the sign of each of the coefficients (the actual value doesn't matter, just the sign) and as we only have even power terms, f(-x) is also + + + just makes it a bit easier to spot sign changes, at least for me. Your mileage may vary, as they say...

OpenStudy (anonymous):

Ohh okay yeah that makes sense. One last quick question if you dont mind then i gotta go to bed!!! :p

OpenStudy (whpalmer4):

Ask quickly :-)

OpenStudy (anonymous):

Using Descartes? rule of signs, how many real roots does 5x^6 + 4x^5 + 3x^4 - 2x^3 + 2x^2 - x -1

OpenStudy (anonymous):

three or one positive real roots, one or zero negative real roots

OpenStudy (anonymous):

?

OpenStudy (whpalmer4):

\[5x^6 + 4x^5 + 3x^4 - 2x^3 + 2x^2 - x -1\] f(x): + + + - + - - 3 positive real roots, or 3-2=1 positive real root f(-x): + - + + + + - 3 negative real roots, or 3-2=1 negative real root total of 6 roots, so we have either 3 positive, 3 negative or 3 positive, 1 negative, 2 complex or 1 positive, 3 negative, 2 complex or 1 positive, 1 negative, 4 complex

OpenStudy (whpalmer4):

turns out to be the final case

OpenStudy (anonymous):

Okay i counted wrong, im tired :p lol. So three or one positive real roots, three or one negative real roots

OpenStudy (whpalmer4):

Yes, that's correct.

OpenStudy (anonymous):

Okay thank you so much! You've been so much help and helped me understand a lot of stuff that i didnt before :D

OpenStudy (whpalmer4):

Glad I could help! Good night, and hopefully you won't see any equations in your dreams :-)

OpenStudy (anonymous):

Hahaha hopefully not :p. Goodnight and thanks again!